# Efficient Active Imitation Learning with Random Network Distillation

### v1

@@ 1 Introduction
Imitation learning has increasingly become a favored approach for learning behaviors in complex environments, offering a compelling alternative to classical scripted behaviors implemented by domain specialists (Schaal, 1999; Hussein et al., 2017). It is particularly well suited in problems where there is not a clear performance measure (or reward). In video games, it is becoming more and more familiar to game developers that frequently address the problem of implementing bots in their games which must play in realistic ways (Harmer et al., 2018; Yadgaroff et al., 2023; Mao et al., 2024). Indeed, since the notion of realism is not well specified, it prevents the use of reinforcement learning-based approaches where a reward signal is mandatory. By observing and replicating human players, these bots are trained to execute complex strategies and actions that are both efficient and human-like.
@@
In response to these limitations, some approaches involve human-in-the-loop training framework, often referred to as active imitation learning (Ross et al., 2011; Judah et al., 2012; Menda et al., 2019; Hoque et al., 2021). Instead of gathering very large amounts of demonstrations, these methods focus on the strategic incorporation of human feedback to selectively guide the learning process, thereby enhancing the efficiency of the training phase. By prioritizing the acquisition of relevant and impactful expert traces, these approaches seek to expedite the agent’s learning curve without relying on expansive datasets, and swiftly improve performance. In the aforementioned video game example, this corresponds to collecting a first set of players traces to bootstrap the agent using classical imitation learning, and then to gather additional traces in particular race settings during training, i.e. requiring a player to control the car in particular and relevant situations.
Figure 1: RND-DAgger overview. (a) The learner’s policy controls the agent until (b) our Random Network Distillation-based out-of-distribution (OOD) measure is triggered. Then (c) the expert takes control until the OOD measure gets lower than the threshold for at least W steps. (d) At last, the current policy takes control of the agent to continue the episode, and can trigger the expert again later.
@@
Our main contributions are threefold: i) We propose a new method called RND-DAgger which is a novel interactive imitation learning approach leveraging state-based out-of-distribution identification through random network distillation. ii) We perform a comparative analysis of RND-DAgger and existing methods on 3 tasks: a robotics scenario and two video-game environments. iii) Throughout these experiments, we demonstrate that RND-DAgger either outperforms or matches existing approaches in terms of final performance while significantly reducing expert burden.
@@ 2 Preliminaries and Related Models
Let us consider a Partially Observable Markov Decision Process described by a state space $\mathcal{S}$, an observation space $\mathcal{O}$ and an action space $\mathcal{A}$. The dynamics of the process is defined by two unknown distributions: i) the observation distribution $P(o_{t}|s_{t})$ with $(o_{t},s_{t})\in\mathcal{O}\times\mathcal{S}$ and ii) the transition function $P(s_{t+1}|a_{t},s_{t})$ where $a_{t}$ lies in $\mathcal{A}$. Note that we do not consider any reward in this setting since the task to solve is not explicit (e.g. driving a car), our objective being typically to imitate human behaviors that may be suboptimal and diverse (e.g. if considering multiple players or experts). An episode $\tau$ is composed of a sequence of observations and actions $\tau=(o_{1},a_{1},...,o_{T_{\#\tau}},a_{T_{\#\tau}})$ where #$\#\tau$ is the size of the episode. Given a policy $\pi(a_{t}|o_{t},a_{t-1},o_{t-1},....,o_{1})$, it is possible to sample an episode by sequentially executing the policy until reaching a stop criterion.
@@
4:$t\leftarrow 0$
5:for $i=0,\ldots,K$ DAgger iterations do
6:     $D_{i}\leftarrow\emptyset$
7:     for $T$ sampling steps do
8:         $s_{t}\sim Env(a_{t-1})$ (Get a new state)
9:         $a_{t}\leftarrow\beta_{i}\pi_{exp}(s_{t})+(1-\beta_{i})\pi_{i}(s_{t})$
10:         $D_{i}\leftarrow D_{i}\cup\{(a_{t},\pi_{exp}(s_{t}))\}$
11:         $t\leftarrow t+1$
12:     end for
13:     Push $D\leftarrow D\cup D_{i}$
14:     Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$
15:end for
16:return $\pi_{K}$
@@
6:$D_{i}\leftarrow\emptyset$
7:for $i=0,\ldots,K$ iterations do
8:     while #$\#D_{i}<T$  do
9:         $s_{t}\sim Env(a_{t-1})$ (Get a new state)
10:         $C_{t}\leftarrow\textsc{Condition}(\pi_{exp},\pi_{i},s_{t})$
11:         if $C_{t}$ then
12:              $a_{t}\leftarrow\pi_{exp}(s_{t})$
13:              $D_{i}\leftarrow D_{i}\cup\{(a_{t},s_{t})\}$
14:              $t\leftarrow t+1$
15:              if not $\textbf{not }C_{t-1}$ then
16:                  $nswitch\leftarrow nswitch+1$
17:              end if
18:         else
19:              $a_{t}\leftarrow\pi_{i}(s_{t})$
20:         end if
21:         $t\leftarrow t+1$
22:     end while
23:     $D\leftarrow D\cup D_{i}$
24:     Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$
25:end for
26:return $\pi_{K}$
@@
To better sample relevant states, few variants of the DAgger algorithm have been proposed. One of the most relevant is Ensemble-DAgger (Menda et al., 2019) which introduces a decision rule to decide if the agent is controlled by the current policy, or by the expert one. Moreover, pairs of state-actions are added to the learning dataset only when the action comes from the expert policy (see Figure 2 - right ). The decision rule is a discrepancy measure that computes the distance between the expert action and the policy action (combined with a disagreement measure between a mixture of experts) as shown in Algorithm 3. This approach is not realistic when considering human experts. Indeed, let us revisit our car driving example, where the goal is to learn an effective driving policy for a video game. Using the DAgger and EnsembleDAgger approaches, the expert (i.e. the player) must continuously play the game during the active imitation learning process, while the car is only partially controlled by the current learned policy. This setup has two significant drawbacks: i) First, the player is placed in an unnatural setting, where they are expected to control a car that they do not fully manage. This can create a disorienting experience, as the player must constantly adapt to actions taken by the policy, disrupting the natural flow of gameplay. ii) Second, the player remains engaged even when the car is exhibiting good behavior, spending time providing feedback that may not significantly contribute to discovering a better policy. This leads to inefficiencies, as the expert’s time is not always optimally utilized, especially in situations where the current policy already performs well.
Algorithm 3 Condition for Ensemble-DAgger
@@
4:if $\sigma^{2}_{nov}>\chi\textbf{ or }\text{measure}>\tau$ then
5:     return True
6:else
7:     return False
8:end if
Algorithm 4 Condition for Lazy-DAgger
@@
$m_{t}\leftarrow f(s_{t})\text{ which estimates }\|\pi_{exp}(s_{t})-\pi_{i}(s_{t})\|^{2}$
3:if $m_{t}>\beta_{H}$ then
4:     return True
5:else
6:     if $\text{measure}<\beta_{R}$ then
7:         return False
8:     else
9:         return True
10:     end if
11:end if
Figure 3: In Ensemble-DAgger, the policy is a mixture, and the measure is based on both a disagreement between the models in this mixture, and a disagreement with the expert action. The computations of these measure require the expert to provide action at each timestep of the process. In Lazy-DAgger, the measure is based on the disagreement between the current policy and the expert one, but to avoid the query of the expert action at each timestep, a classifier $f$ is trained to predict if the discrepency measure $\|\pi_{exp}(s_{t})-\pi_{i}(s_{t})\|^{2}$ is above a given threhsold $\beta_{H}$ or not.
@@ 3 Random Network Distillation-based DAgger
6:$D_{0}\leftarrow\emptyset$
7:for $i=0,\ldots,K$ DAgger iterations do
8:     while #$\#D_{i}<T$  do
9:         $s_{t}\sim Env(a_{t-1})$ (Get a new state)
10:         $m_{t}\leftarrow\|f_{targ}(s_{t})-f_{pred}(s_{t})\|^{2}$
11:         if $m_{t}>\lambda$ or $w<W$ then
12:              if $m_{t}\leq\lambda$ then (Minimal demo time)
13:                  $w\leftarrow w+1$
14:              else
15:                  $w\leftarrow 0$
16:              end if
17:              $a_{t}\leftarrow\pi_{exp}(s_{t})$
18:              $D_{i}\leftarrow D_{i}\cup\{(a_{t},s_{t})\}$
19:              if $m_{t-1}\leq\lambda$ then
20:                  $nswitch\leftarrow nswitch+1$
21:              end if
22:         else
23:              $w\leftarrow 0$
24:              $a_{t}\leftarrow\pi_{i}(s_{t})$
25:         end if
26:         $t\leftarrow t+1$
27:     end while
28:     Push $D\leftarrow D\cup D_{i}$
29:     Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$
30:     Train $\text{Train }f_{pred}\text{ on }D\text{ to predict }f_{targ}$
31:end for
32:return $\pi_{K}$
@@
Lazy-DAgger replaces the traditional action-based discrepancy measure between expert and policy actions with a classifier-based approximation — see Figure 4. This approach allows the algorithm to predict when expert intervention is necessary without requiring an expert action at every timestep, thereby significantly reducing the expert burden compared to methods like Ensemble-DAgger. However, this method has a notable drawback: similar to Ensemble-DAgger, it relies on comparing the actions of the expert and the current policy for a given state. While this approach works well when the expert is optimal and acts deterministically, it becomes problematic when dealing with humans or imperfect experts. Human experts often exhibit variability and may choose different actions for the same state depending on context or personal preference. This results in a noisy and unreliable measure of discrepancy, causing the method to trigger unnecessary interventions or overlook critical corrections, ultimately reducing its effectiveness.
We propose the RND-DAgger algorithm, which builds upon Random Network Distillation (RND) (Burda et al., 2018). The core assumption of RND-DAgger is that expert feedback is only necessary when the agent encounters out-of-distribution (OOD) states—states that are not well represented in the training set and where the policy is more likely to fail. This is crucial because when an agent operates in OOD states, it faces a higher risk of taking suboptimal actions that could hinder learning or lead to unsafe outcomes. Once the agent returns to familiar, in-distribution states, the expert’s feedback is no longer required, and the intervention is ended. Our method differs fundamentally from existing approaches like Ensemble-DAgger and Lazy-DAgger, which rely on action-based discrepancy measures to determine when to intervene. Thus, our approach remains robust to variations in expert behavior and ensures that interventions only occur when the agent genuinely faces unfamiliar and potentially risky states.
@@
To measure if a state is OOD, RND-DAgger relies on the Random Network Distillation technique, which is a classical approach developed initially for the problem of exploration in Reinforcement Learning to detect new states to explore (Burda et al., 2018). The principle of RND is to use a randomly initialized neural network as a fixed target $f_{targ}$ and train a second neural network $f_{pred}$ (the predictor) to approximate the output of the target network. As the predictor network improves over time, the error between the predictor and the target network decreases for familiar states (in-distribution), but remains high for unseen or out-of-distribution states. This prediction error thus serves as a measure of novelty, allowing the agent to recognize when it has encountered a new or unfamiliar state. In RND-DAgger, the measure is used to decide when to trigger the expert, and when to trigger back the current policy, alternating phases where the expert is controlling the agent and when the agent is controlled by the current policy. Note that, contrarily to the described baselines, this measure does not involve the expert action, thus avoiding the need to have an expert acting at each timestep.
@@
Just relying on the OOD measure can lead to a behavior where the expert is asked too often for short demonstrations (see appendix Figure 13(d)). To account for this expert burden, we introduce the notion of minimal expert time. It is defined as $W$ the number of consecutive "in-distribution" frames to switch control back from the expert to the autonomous policy (e.g. then number of frames where $m_{t}<\|f_{targ}(s_{t})-f_{pred}(s_{t})\|^{2}$). This ensures that the expert will provide a minimal of information at each intervention, thus demonstrating fewer sequences but of longer length.
@@ 4 Experiments
Our first environment is HalfCheetah which is a classical reinforcement learning environment [^1] https://github.com/araffin/pybullet_envs_gymnasium where the objective is to learn a running strategy for the agent. The goal of the agent is to locomote as fast as possible. The HalfCheetah is controlled by applying motor torques, and the agent manipulates a 6-degree-of-freedom (6-DOF) motor joint vector. The observation space consists of 18 values, which include the position and velocity of the agent’s body, the angles and angular velocities of its six joints, as well as the angle of the center of mass. We also propose two new environments developed for video game research [^2] The code of these environments that have been made using the Godot engine will be released upon acceptance. RaceCar (see Figure 5) features a physics-based car controller that must complete a single lap on a given track. After each lap, the car is reset to a random position on the starting line. The track presents several challenges, including speed bumps, a ramp in front of a pillar, and sharp 90-degree turns, making the optimal driving behavior non-trivial. Additionally, crossing the ramp and red sloped walls provides a speed boost, adding complexity to the strategy. The car is controlled using four discrete actions: forward, backward, left, and right. The observation space consists of 22 dimensions, including the car’s 3D position, linear velocity (3D), angular velocity (3D), rotation (encoded as the cosine and sine of the angle), and data from seven raycasts to detect obstacles. At last, the 3D Maze environment allows us to study our strategy in goal-conditioned navigation scenarios. A classical human-like character is spawned at a random location within the maze and tasked with reaching a randomly assigned goal. The agent is controlled using four movement actions: walk forward, walk backward, strafe left, and strafe right, as well as a one-dimensional rotation action to change its facing direction.The observation space consists of the agent’s absolute position (3D), the goal’s position (3D), and 131 raycasts, which provide a low-resolution depth map of the surroundings. Since the objective in this environment is to navigate to any goal location, we replace the traditional Behavioral Cloning (BC) approach with a Goal-Conditioned Behavioral Cloning (GCBC) approach (Ding et al., 2019b) as the base learning algorithm. This adaptation ensures that the agent learns not just to imitate expert trajectories, but also to generalize its navigation strategies based on varying goal positions within the maze.
@@
Figure 5: Illustration of the three different environments used in our experiments. The objective is to learn a good driving/navigating/walking policy while minimizing the interventions of the expert.
@@
We compare our approach RND-DAgger to multiple approaches from the active imitation learning literature: DAgger Ross et al. (2011), a classical approach which queries the supervisor for an action in every state that the learner visits. Ensemble-DAgger Menda et al. (2019), which propose an automated approach to measure OOD and trigger expert demonstration. Ensemble-DAgger relies on two metrics: A disagreement measure defined as the correlation between an ensemble of policies, and a discrepancy measure computing the difference between the bot action and the expert action at every step. Lazy-DAgger uses the same discrepancy measure as Ensemble-DAgger and reduce bot-expert context switching by using an additional threshold parameter to define a hysteresis band. Note that, we rely on a simplified version proposed in (Hoque et al., 2021) that has been shown to be more efficient and generates less context switches than the original version. This variant discards the classifier for the benefit of the true discrepancy measure. Human-Gated DAgger (HG-DAgger)(Kelly et al., 2019) involves a human supervisor monitoring the agent’s behavior in real-time and deciding when to take control. While similar in essence to our proposed RND-DAgger, HG-DAgger relies entirely on human judgment to identify when intervention is necessary, rather than using an automatic OOD measure. This requires the expert to continuously observe the agent’s actions at every timestep, leading to a higher cognitive load and less efficient use of expert time. Finally, we also consider Behavioral Cloning (BC) as the classical baseline, which does not include any active learning mechanism.

### v2

@@ 1 Introduction
Imitation learning has increasingly become a favored approach for learning behaviors in complex environments, offering a compelling alternative to classical scripted behaviors implemented by domain specialists (Schaal, 1999; Hussein et al., 2017). It is particularly well suited in problems where there is not a clear performance measure (or reward). In video games, it is becoming more and more familiar to game developers that frequently address the problem of implementing bots in their games which must play in realistic ways (Harmer et al., 2018; Yadgaroff et al., 2024; Mao et al., 2024). Indeed, since the notion of realism is not well specified, it prevents the use of reinforcement learning-based approaches where a reward signal is mandatory. By observing and replicating human players, these bots are trained to execute complex strategies and actions that are both efficient and human-like.
@@
In response to these limitations, some approaches involve human-in-the-loop training framework, often referred to as active imitation learning (Ross et al., 2011; Judah et al., 2012; Menda et al., 2019; Hoque et al., 2021). Instead of gathering very large amounts of demonstrations, these methods focus on the strategic incorporation of human feedback to selectively guide the learning process, thereby enhancing the efficiency of the training phase. By prioritizing the acquisition of relevant and impactful expert traces, these approaches seek to expedite the agent’s learning curve without relying on expansive datasets, and swiftly improve performance. In the aforementioned video game example, this corresponds to collecting a first set of player traces to bootstrap the agent using classical imitation learning, and then to gather additional traces in particular race settings during training, i.e. requiring a player to control the car in particular and relevant situations.
Figure 1: RND-DAgger overview. (a) The learner’s policy controls the agent until (b) our Random Network Distillation-based out-of-distribution (OOD) measure is triggered. Then (c) the expert takes control until the OOD measure gets lower than the threshold for at least W steps. (d) Finally, the current policy takes control of the agent to continue the episode, and can trigger the expert again later.
@@
Our main contributions are threefold: i) We propose a new method called RND-DAgger, a novel interactive imitation learning approach leveraging state-based out-of-distribution identification through random network distillation. ii) We perform a comparative analysis of RND-DAgger and existing methods on 3 tasks: a robotics scenario and two video-game environments. iii) Throughout these experiments, we demonstrate that RND-DAgger either outperforms or matches existing approaches in terms of final performance while significantly reducing expert burden.
@@ 2 Preliminaries and Related Models
Let us consider a Markov Decision Process described by a state space $\mathcal{S}$ and an action space $\mathcal{A}$. The dynamics of the process are defined by an unknown transition function $P(s_{t+1}|a_{t},s_{t})$, where $a_{t}\in\mathcal{A}$ and $s_{t}\in\mathcal{S}$. Note that we do not consider any reward in this setting since the task to solve is not explicit (e.g., driving a car); our objective is instead to imitate human behaviors that may be suboptimal and diverse (e.g., considering multiple players or experts). An episode $\tau$ consists of a sequence of states and actions, $\tau=(s_{1},a_{1},\ldots,s_{T},a_{T})$, where $T$ is the length of the episode. Given a policy $\pi(a_{t}|s_{t})$, it is possible to sample an episode by sequentially executing the policy until reaching a stop criterion.
@@
[…]
6:$s_{t}\sim Env(a_{t-1})$ (Get a new state)
7:$a_{t}\leftarrow\beta_{i}\pi_{exp}(s_{t})+(1-\beta_{i})\pi_{i}(s_{t})$
8:$D_{i}\leftarrow D_{i}\cup\{(a_{t},\pi_{exp}(s_{t}))\}$
9:$t\leftarrow t+1$EndFor
10:Push $D\leftarrow D\cup D_{i}$
11:Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$EndFor
12:return $\pi_{K}$
@@
[…]
7:$s_{t}\sim Env(a_{t-1})$ (Get a new state)
[…]
9:$a_{t}\leftarrow\pi_{exp}(s_{t})$
10:$D_{i}\leftarrow D_{i}\cup\{(a_{t},s_{t})\}$
11:$t\leftarrow t+1$Ifnot $\textbf{not }C_{t-1}$
[…]
13:$a_{t}\leftarrow\pi_{i}(s_{t})$EndIf
14:$t\leftarrow t+1$EndWhile
15:$D\leftarrow D\cup D_{i}$
16:Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$EndFor
17:return $\pi_{K}$
@@
To better sample relevant states, few variants of the DAgger algorithm have been proposed. One of the most relevant is Ensemble-DAgger (Menda et al., 2019) which introduces a decision rule to decide if the agent is controlled by the current policy, or by the expert one. Moreover, pairs of state-actions are added to the learning dataset only when the action comes from the expert policy (see Figure 2 - right ). The decision rule is a discrepancy measure that computes the distance between the expert action and the policy action (combined with a disagreement measure between a mixture of experts) as shown in Algorithm 3. This approach is not realistic when considering human experts. Indeed, let us revisit our car driving example, where the goal is to learn an effective driving policy for a video game. Using the DAgger and Ensemble-DAgger approaches, the expert (i.e. the player) must continuously play the game during the active imitation learning process, while the car is only partially controlled by the current learned policy. This setup has two significant drawbacks: i) First, the player is placed in an unnatural setting, where they are expected to control a car that they do not fully manage. This can create a disorienting experience, as the player must constantly adapt to actions taken by the policy, disrupting the natural flow of gameplay. ii) Second, the player remains engaged even when the car is exhibiting good behavior, spending time providing feedback that may not significantly contribute to discovering a better policy. This leads to inefficiencies, as the expert’s time is not always optimally utilized, especially in situations where the current policy already performs well.
Algorithm 3 Ensemble-DAgger’s Condition
@@
[…]
Algorithm 4 Lazy-DAgger’s Condition
@@
[…]
Figure 3: In Ensemble-DAgger, the policy is a mixture, and the measure is based on both a disagreement between the models in this mixture, and a disagreement with the expert action. The computations of these measure require the expert to provide action at each timestep of the process. In Lazy-DAgger, the measure is based on the disagreement between the current policy and the expert one, but to avoid the query of the expert action at each timestep, a classifier $f$ is trained to predict if the discrepancy measure $\|\pi_{exp}(s_{t})-\pi_{i}(s_{t})\|^{2}$ is above a given threshold $\beta_{H}$ or not.
@@ 3 Random Network Distillation-based DAgger
[…]
7:$s_{t}\sim Env(a_{t-1})$ (Get a new state)
[…]
9:$w\leftarrow w+1$Else
10:$w\leftarrow 0$EndIf
11:$a_{t}\leftarrow\pi_{exp}(s_{t})$
[…]
14:$w\leftarrow 0$
15:$a_{t}\leftarrow\pi_{i}(s_{t})$EndIf
16:$t\leftarrow t+1$EndWhile
17:Push $D\leftarrow D\cup D_{i}$
18:Train $\text{Train }\pi_{i+1}\text{ on }D\text{ using BC}$
19:Train $\text{Train }f_{pred}\text{ on }D\text{ to predict }f_{targ}$EndFor
20:return $\pi_{K}$
@@
Lazy-DAgger replaces the traditional action-based discrepancy measure between expert and policy actions with a classifier-based approximation — see Figure 4. This approach allows the algorithm to predict when expert intervention is necessary without requiring an expert action at every timestep, thereby significantly reducing the expert burden compared to methods like Ensemble-DAgger. However, this method has a notable drawback: similar to Ensemble-DAgger, it relies on comparing the actions of the expert and the current policy for a given state. While this approach works well when the expert is optimal and acts deterministically, it becomes problematic when dealing with humans or imperfect experts. Human experts often exhibit variability and may choose different actions for the same state depending on context or personal preference. This results in a noisy and unreliable measure of discrepancy. To address this issue, one could rely on a state-based discrepancy measure that would focus on meaningful divergences compared to an action-based one.
We propose the RND-DAgger algorithm, which builds upon Random Network Distillation (RND) (Burda et al., 2019). The core assumption of RND-DAgger is that expert feedback is only necessary when the agent encounters out-of-distribution (OOD) states—states that are not well represented in the training set and where the policy is more likely to fail. This is crucial because when an agent operates in OOD states, it faces a higher risk of taking suboptimal actions that could hinder learning or lead to unsafe outcomes. Once the agent returns to familiar, in-distribution states, the expert’s feedback is no longer required, and the intervention is ended. Our method differs fundamentally from existing approaches like Ensemble-DAgger and Lazy-DAgger, which rely on action-based discrepancy measures to determine when to intervene. As such, our approach remains robust to variations in expert behavior, ensuring that interventions only occur when the agent genuinely faces unfamiliar and potentially risky states.
@@
To measure if a state is OOD, RND-DAgger relies on the Random Network Distillation technique, which is a classical approach developed initially for the problem of exploration in Reinforcement Learning to detect new states to explore (Burda et al., 2019). The principle of RND is to use a randomly initialized neural network as a fixed target $f_{targ}$ and train a second neural network $f_{pred}$ (the predictor) to approximate the output of the target network. As the predictor network improves over time, the error between the predictor and the target network decreases for familiar states (in-distribution), but remains high for unseen or out-of-distribution states. This prediction error thus serves as a measure of novelty, allowing the agent to recognize when it has encountered a new or unfamiliar state. In RND-DAgger, the measure is used to decide when to trigger the expert, and when to trigger back the current policy, alternating phases where the expert is controlling the agent and when the agent is controlled by the current policy. Note that, contrary to the described baselines, this measure does not involve the expert action, thus avoiding the need to have an expert acting at each timestep.
@@
Relying exclusively on the OOD measure can lead to a behavior where the expert is asked too often for short demonstrations (see appendix Figure C.2). To account for this limitation, we introduce the notion of minimal expert time (MET). It is defined as $W$ the number of consecutive "in-distribution" frames required to switch control back from the expert to the autonomous policy (i.e. the number of frames where $m_{t}<\|f_{targ}(s_{t})-f_{pred}(s_{t})\|^{2}$). This mechanism ensures that the expert will provide a minimal amount of information at each intervention, resulting in fewer, but longer, demonstration sequences.
@@ 4 Experiments
Our first environment is HalfCheetah which is a classical reinforcement learning environment [^1] https://github.com/araffin/pybullet_envs_gymnasium where the objective is to learn a running strategy for the agent. The goal of the agent is to locomote as fast as possible. The HalfCheetah is controlled by applying motor torques, and the agent manipulates a 6-degree-of-freedom (6-DOF) motor joint vector. The observation space consists of 18 values, which include the position and velocity of the agent’s body, the angles and angular velocities of its six joints, as well as the angle of the center of mass. We also propose and open-source two new environments developed for video game research . RaceCar (see Figure 5) features a physics-based car controller that must complete a single lap on a given track. After each lap, the car is reset to a random position on the starting line. The track presents several challenges, including speed bumps, a ramp in front of a pillar, and sharp 90-degree turns, making the optimal driving behavior non-trivial. Additionally, crossing the ramp and red sloped walls provides a speed boost, adding complexity to the strategy. The car is controlled using four discrete actions: forward, backward, left, and right. The observation space consists of 22 dimensions, including the car’s 3D position, linear velocity (3D), angular velocity (3D), rotation (encoded as the cosine and sine of the angle), and data from seven raycasts to detect obstacles. Finally, the 3D Maze environment allows us to study our strategy in goal-conditioned navigation scenarios. A classical human-like character is spawned at a random location within the maze and tasked with reaching a randomly assigned goal. The agent is controlled using four movement actions: walk forward, walk backward, strafe left, and strafe right, as well as a one-dimensional rotation action to change its facing direction.The observation space consists of the agent’s absolute position (3D), the goal’s position (3D), and 131 raycasts, which provide a low-resolution depth map of the surroundings. Since the objective in this environment is to navigate to any goal location, we replace the traditional Behavioral Cloning (BC) approach with a Goal-Conditioned Behavioral Cloning (GCBC) approach (Ding et al., 2019b) as the base learning algorithm. This adaptation ensures that the agent learns not just to imitate expert trajectories, but also to generalize its navigation strategies based on varying goal positions within the maze.
@@
Figure 5: Illustration of the three different environments used in our experiments. The objective is to learn a good driving, navigating, and walking policy while minimizing expert interventions.
@@
We compare our approach RND-DAgger to multiple approaches from the active imitation learning literature: DAgger Ross et al. (2011), a classical approach which queries the supervisor for an action in every state that the learner visits. Ensemble-DAgger Menda et al. (2019), which propose an automated approach to measure OOD and trigger expert demonstration. Ensemble-DAgger relies on two metrics: A disagreement measure defined as the correlation between an ensemble of policies, and a discrepancy measure computing the difference between the bot action and the expert action at every step. Lazy-DAgger uses the same discrepancy measure as Ensemble-DAgger and reduces bot-expert context switching by using an additional threshold parameter to define a hysteresis band. Note that, we rely on a simplified version proposed in (Hoque et al., 2021) that has been shown to be more efficient and generates less context switches than the original version. This variant discards the classifier for the benefit of the true discrepancy measure. Human-Gated DAgger (HG-DAgger) from Kelly et al. (2019) involves a human supervisor monitoring the agent’s behavior in real-time and deciding when to take control. While similar in essence to our proposed RND-DAgger, HG-DAgger relies entirely on human judgment to identify when intervention is necessary, rather than using an automatic OOD measure. This requires the expert to continuously observe the agent’s actions at every timestep, leading to a higher cognitive load and less efficient use of expert time. Finally, we also consider Behavioral Cloning (BC) as the classical baseline, which does not include any active learning mechanism.